Suppose the following information describes the economy of a hypothetical country::

 

C        = 200 + 0.9Yd

I         = 100

G        = 250

X        = 200

Z         = 50 + 0.12Y

t       = 20%

 

Using the multiplier approach, calculate the equilibrium income.

 

  • A. 1925
  • B. 1750
  • C. 2950
  • D. 2000
1
Expert's answer
2021-11-14T17:37:17-0500

Solution:

The correct answer is B. 1,750

 

Equilibrium income (Y): AD = AS: Y = Z

Y = C + I + G + X

C = 200 + 0.9Yd = 200 + 0.9(Y – T) = 200 + 0.9(Y – 0.2Y) = 200 + 0.9Y – 0.18Y

Y + 50 + 0.12Y = 200 + 0.9Y – 0.18Y + 100 + 250 + 200

1.12Y + 0.18Y – 0.9Y = 750 - 50

0.4Y = 700

Y = 1750

Equilibrium income = 1,750